# Direct $N$-body simulations of rotating and extremely massive Population III star clusters

### 31 Mar 2026 v1

@@ 4.3.1 Structural parameter evolution
  • Phase IV: from core-collapse to 500 Myr (the simulations’ end). The Lagrangian radii evolve self-similarly, as shown in 8, driven by binary heating in the core which causes global expansion. Models with larger $\omega_{0}$ expand more by 500 Myr because they undergo earlier core collapse; their subsequent $r_{\mathrm{Lagr}}$ evolution is otherwise similar. The clusters become fully mass segregated ($M_{\mathrm{av}}$ $\approx 200\penalty 10000\ \mathrm{M}_{\odot}{}$ centrally, $\approx 30\penalty 10000\ \mathrm{M}_{\odot}{}$ in the halo). The same long-term trends appear in Fig. 3; longer runs may amplify differences. These results are consistent with energy-flow theorems across the half-mass radius $r_{\mathrm{h}}$ (Hénon, 1975; Breen & Heggie, 2013).
@@
The gravothermal-gravogyro catastrophe is therefore present in all our Pop III cluster models, but its timescale depends strongly on the initial rotation. GRK have little effect on the global structural evolution.
@@ 4.3.2 Angular momentum transport
We explore whether and how angular-momentum redistribution depends on stellar evolution, GRK, and initial bulk rotation. To track the gravothermal-gravogyro catastrophe, we divide the ZAMS particles into four classes: low-mass core-collapse supernovae (lCCSNe): $8.0\leq m_{\mathrm{ZAMS}}<85.0\penalty 10000\ \mathrm{M}_{\odot}$, pulsational pair-instability supernovae (PPISNe): $85.0\leq m_{\mathrm{ZAMS}}<120.0\penalty 10000\ \mathrm{M}_{\odot}$, pair-instability supernovae (PISNe): $120.0\leq m_{\mathrm{ZAMS}}<265.0\penalty 10000\ \mathrm{M}_{\odot}$, and, high-mass core-collapse supernovae (hCCSNe): $265.0\leq m_{\mathrm{ZAMS}}<300.0\penalty 10000\ \mathrm{M}_{\odot}$, where $m_{\mathrm{ZAMS}}$ is the ZAMS stellar mass of a single star [^6] A primordial binary may have members of two different mass groups.. We track particles from each group over the full evolution and compute their total angular momentum squared $L_{\mathrm{group}}^{2}$ as Kamlah et al. (2022b) normalized by $L^{2}_{\omega_{0}=0.6,t=0}$ for the (No)K0.6 model. Fig. 11 shows these quantities together with the group masses $M_{\mathrm{group}}$.
@@
Phase II: $\sim$2-3 Myr. Rapid stellar evolution drives a sharp fall in $M_{\mathrm{group}}$, with higher-mass groups declining earlier owing to their shorter evolution timescales. PISNe remove both mass and a large fraction of angular momentum: $M_{\mathrm{PISNe}}$ carried the most angular momentum before these events and is effectively depleted afterwards, causing a large loss of rotational kinetic energy.
Phase III: $\sim$3 Myr to core collapse (around $\sim$50-80 Myr depending on $\omega_{0}$. See Sect. 4.3.1; Fig. 8). The quantity $L_{\mathrm{hCCSNe}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$ initially falls after phase II but, in the (No)K1.2 and (No)K1.8 models, it instead rises sharply and peaks at $\sim$10-20 Myr. After the peak, the quantity of (No)K1.2 shows a gradual decline while that of (No)K1.8 drops rapidly. Other mass groups follow the same pattern except $M_{\mathrm{PISNe}}$. In (No)K1.8 the quantity attains a local minimum at core collapse, when the cluster is essentially fully mass segregated (Fig. 8).
@@
Compared to rotating Pop II clusters (Kamlah et al., 2022b), the detailed transport history differs because of the IMF and stellar evolution, but the gravothermal-gravogyro catastrophe remains visible in both cases. GRK again produce no strong effect.
@@ 4.4 Collision and coalescence events
We define a “coalescence” as the merger of a nearly circular binary, while “collisions” are a direct physical impact either on a hyperbolic or highly eccentric orbit; this is also how the Nbody6++GPU code separates them. Under our initial conditions and Pop III stellar evolution prescriptions, we find collisions of BHBH, MSBH and MSMS, while coalescences may involve MS, HG, CHeB, ShHeB, HeMS, NS and BH progenitors. Fig. 12 shows the product masses $m_{\mathrm{coal}}$ and $m_{\mathrm{coll}}$ versus $r_{\mathrm{dens}}$. For coalescence, only MS, CHeB, ShHeB and BH products appear, consistent with the merger-product treatment of Hurley et al. (2002) and Kamlah et al. (2022a). Event totals are given in Table 4.
@@
Our suite contains one realization per setup ($N=1$ per model ID). For integer event counts $N$ (specific coalescence/collision types, compact-object merger channels), the counting uncertainty is well approximated by Poisson statistics, with a reference $1\sigma$ uncertainty $\sigma_{N}\simeq\sqrt{N}$ (relative uncertainty $\sigma_{N}/N\simeq 1/\sqrt{N}$). Consequently, differences of a few events are not, by themselves, strong evidence for a systematic dependence on $\omega_{0}$ when NN is in the single-digit to few-tens regime [^7] By contrast, trends inferred from continuous structural quantities (e.g. Fig. 3, Fig. 8, and the angular-momentum transport in Fig. 11) and from large-number counts (e.g. total escaper numbers; Table 5) are dynamically more robust in our dataset..
@@
The coalescence events show no marked spatial differences: most occur within the central 2.5 pc, i.e. inside $r_{h}$ during the first $\sim$40 Myr (Fig. 3). They are mostly completed by 4-5 Myr (pre-core-collapse, Fig. 9). Many of these early mergers populate the PISNe mass gap, indicating that stellar mergers in early star cluster evolution can seed BHs in and around the gap (see, e.g., Costa et al., 2022; Ballone et al., 2022, for recent hydrodynamical simulations of stellar collisions on this topic). With our flat IMF, Pop III stellar evolution, and absent wind mass loss, producing BHs above the PISNe gap is easier here than in typical young massive clusters (e.g., Rizzuto et al., 2022).
@@
Most coalescences produce CHeB stars and the most massive products, typically occurring within a few Myr in all runs. Faster initial rotation is consistent with more CHeB coalescences (Fig. 9), but Poisson uncertainties and covariances between channels in a single realization make this indicative. The numbers of BH, ShHeB and MS coalescences are similar across simulations. These mergers produce stars (mainly CHeB) with masses inside and above the pair-instability gap; some IMBHs end up in the pair-instability gap after accreting stellar material and grow further.
@@
MSMS, BHMS and BHBH collisions offer useful diagnostics. Collision-produced IMBH masses are broadly similar across models (Fig. 12), except NoK1.2 (IMBH masses $\approx$ 1000 $\mathrm{M}_{\odot}$) which is not seen in K1.2. Fig. 9 suggests more MSMS and BHBH collisions at higher initial rotation, but given the small absolute counts the implied trend should be read as suggestive and may be affected by stochastic fluctuations. Notably, K models show multi-generation IMBH growth.
@@
Fig. 13 shows the primary vs. secondary masses for MSMS, BHMS and BHBH collisions across all runs; the distributions are similar regardless of rotation or GRK. Contrary to theoretical expectations (e.g. Morawski et al., 2018, 2019) [^8] From these theories, we would expect that mergers have very low GRK velocity, which would make them promising candidates for growing seed BHs for galactic nuclei., Fig. 14 demonstrates that most GRK velocities $v_{\mathrm{GRk}}$ exceed cluster escape speeds in the core, so merger remnants are often ejected even for $q\approx 1$. Hence, IMBHs of $1000\mathrm{M}_{\odot}{}$ are unlikely to form via hierarchical merging in Pop III clusters unless additional processes retain the remnants [^9] A surrounding dark matter halo is unlikely to help to retain merger remnants. Even assuming a NFW big-halo of 100 times of the cluster total mass with 10 times of the cluster tidal radius will only contribute to a few $\mathrm{km\penalty 10000\ s}^{-1}$ of the escape speed in the core., such as strong gas inflow after cluster formation causing extreme contraction of the cluster (Kroupa et al., 2020).
@@
To summarize, with increasing initial rotation $\omega_{0}$, coalescences occur earlier and more frequently, and there is a tentative rise in MSMS and BHBH collision counts limited by low-number statistics. Enabling GRK generates kick velocities that frequently exceed the cluster escape speed, favouring remnant ejection and thus suppressing collision-built IMBHs, even though multi-generation IMBH growth can still occur.
@@ 4.5 Escaper stars
Figure 4: Plots showing cumulative counts of escaping single stars $n_{\mathrm{esc,singles}}$, their total mass $M_{\mathrm{esc,singles}}$, and counts of escaping single MS, CHeB, ShHeB, NS and BH stars ($n_{\mathrm{escMS,s}}$, $n_{\mathrm{escCHeB,s}}$, $n_{\mathrm{escShHeB,s}}$, $n_{\mathrm{escNS,s}}$, $n_{\mathrm{escBH,s}}$) for all eight simulations. K: solid; NoK: dash-dotted.
Fig. 4 (single escapers), Fig. 10 (binary escapers) and Table 5 show that higher initial bulk rotation $\omega_{0}$ is associated with earlier and stronger tidal mass loss in both NoK and K models, most clearly for (No)K1.8 and (No)K1.2. The (No)K0.6 models show a non-monotonic behaviour, having $n_{\mathrm{esc,singles}}$, $n_{\mathrm{esc,binarymembers}}$ and $M_{\mathrm{esc,singles}}$ comparable to or below the non-rotating case. A plausible physical interpretation is that two effects compete: (i) rotating King initial conditions are more compact than their non-rotating counterparts (Sect. 1), which can reduce early tidal overflow, while (ii) stronger rotation accelerates early angular-momentum transport and structural readjustment (Sect. 4.3.1), which enhances mass loss once the system expands and interacts with the tidal boundary. In this framework, $\omega_{0}=0.6$ can fall in an intermediate regime where the increased initial compactness dominates the net mass loss, whereas for larger $\omega_{0}$ the rotation-driven early evolution dominates. This tidal field mass loss modifies cluster time-scales and links directly to the structural and angular-momentum evolution discussed in Sects. 4.3.1 and 4.3.2.
@@
We next consider evolved phases, starting with escaping single CHeBs ($n_{\mathrm{escCHeB,s}}$, Fig. 4 panel four). Their behaviour resembles that of escaping MSs, though differences between (No)K1.2 and (No)K1.8 are reduced. Many escaping MSs rapidly become CHeBs, and likewise many CHeBs shortly become ShHeBs (see Fig. 5). After formation, most will remain in the ShHeB phase for comparatively long.
@@
For escaping binary components excluding compact objects, a few MS escapers (counted by $n_{\mathrm{escMS,b}}$) occur only in the (No)K1.2 and (No)K1.8 runs; they are mostly from primordial binaries, low-mass, and unlikely to form IMBHs. CHeB binary escapers ($n_{\mathrm{escBH,b}}$) are rarer still. ShHeB binary losses ($n_{\mathrm{escShHeB,b}}$) are roughly twice as large and occur earlier in the $\omega_{0}=1.2,1.8$ models than in the $\omega_{0}=0.0,0.6$ models. Many late-time escaping BHBH binaries contain two IMBHs and thus cannot fuel hierarchical in-cluster growth. The stronger-rotation models are therefore consistent with a reduced retained BH progenitor/seed population at late times.
Overall, enabling GRK produces the charted differences: with GRK enabled the (No)K1.8 models shows more CHeB single-star losses and fewer MS binary losses, while the (No)K1.2 models shows the opposite trend. We again note that this is limited by small-number statistics.
@@ 5 Summary and conclusions
We present a set of a few pilot direct N-body simulations that probe how bulk rotation and ultra-low metallicity with a top-heavy IMF shape Pop III cluster dynamics. Rotation follows generalized rotating King models (Einsel & Spurzem, 1999) with rigid core rotation and differentially decreasing near the half-mass radius and into the halo. Kamlah et al. (2022b) analysed the general evolution and angular momentum transport in such models of Pop I star clusters; we find generally analogous dynamics, such as the gravogyro and gravothermal collapse, enhanced by the top-heavy IMF (8.08.0300.0300.0 $\mathrm{M}_{\odot}$, Lazar & Bromm, 2022). We study the impact of switching on and off general relativistic merger recoil kicks (GRK) (Arca Sedda et al., 2024), and how the newly implemented stellar evolution for extremely metal-poor stars (EMP, here used $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, Tanikawa et al., 2020) affects the formation and evolution of compact objects. While the mass function of Pop III star clusters is highly uncertain (see Sect. 1 and Stacy et al., 2016; Liu et al., 2021a), our pilot simulations model very massive Pop III star clusters (Table 1) to cover an unexplored parameter range and also because of stronger and faster dynamical effects, such as massive object formation, at high density (Vergara et al., 2025b). Direct N-body simulations using the PeTar code (Wang et al., 2020) for Pop III star clusters (Liu et al., 2024a, b, 2025) have used smaller cluster masses and particle numbers, requiring many independent realizations to average $N$ fluctuations. On the contrary, we use very large particle numbers ($1.01\times 10^{5}$ initially), which suffice to probe global dynamical effects of rotation and gravogyro-gravothermal collapse, but still provide relatively small database for rare binaries and escapers (e.g. those involving black holes).
We ran the models for 500 Myr, using one non-rotating and three rotating King models (Einsel & Spurzem 1999, $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$), each with and without GR merger recoil kicks.
@@
Primordial binary statistics are another major uncertainty for Pop III clusters. In this work we fix the initial binary fraction to $f_{\mathrm{b}}=1\%$ (Table 1), which should be regarded as a low-$f_{\mathrm{b}}$ limiting case rather than a best-estimate value. This choice suppresses IMBH formation pathways tied to primordial binary evolution (in particular Type-I and Type-III; Sect. 4.1) and, by construction, increases the relative importance of dynamical formation channels in our event budget. If $f_{\mathrm{b}}$ were higher, we would expect (i) a larger contribution of mergers from primordial binary channel, and (ii) stronger binary heating which can delay gravothermal collapse and modify the balance between in-cluster hierarchical growth and dynamical ejections. The net change in the total IMBH production efficiency is therefore not straightforward to predict without additional simulations, but the dominance of dynamical channels found here should be interpreted as an upper bound. A dedicated scan in $f_{\mathrm{b}}$ is left for future work.
@@
The global dynamical evolution can be divided into four phases (Figs. 3, 8, 11):
@@
  • Phase I (0-2 Myr): global mass ($M_{\mathrm{cl}}$) and tidal radius ($r_{\mathrm{tid}}$) remain essentially constant because mass loss is negligible; however initial bulk rotation ($\omega_{0}$) controls the structural response. Rapid angular-momentum transport occurs earlier and stronger for larger $\omega_{0}$, producing a brief, synchronous bump in the Lagrangian radii near $\sim$0.1 Myr, inner-radius contraction and outer-radius expansion, and temporary rises then falls of $r_{\mathrm{h}}$, $r_{\mathrm{c}}$ and $m_{\mathrm{c}}$ in the faster-rotating models. Low/zero rotation models shows little change in $r_{\mathrm{h}}$ over this phase. Also, $M_{\mathrm{group}}$ evolves similarly across mass groups while $L_{\mathrm{group}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$ is larger for higher $\omega_{0}$, reflecting early outward angular-momentum migration.
@@
  • Phase II ($\sim$2-3 Myr): rapid stellar-evolution mass loss (notably PISNe and compact-object formation) causes a sharp drop in total mass and $r_{\mathrm{tid}}$, and produces a rapid increase in $r_{\mathrm{h}}$ and $r_{\mathrm{c}}$. PISNe both remove mass and carry away a large fraction of the total angular momentum, while escapers (natal/GR kicks) and dynamical encounters add to mass loss and alter kinematics through fallback-dependent kicks. Measures of central segregation ($M_{\mathrm{av}}$ within small $r_{\mathrm{Lagr}}$) show similar qualitative behaviour across $\omega_{0}$ at this stage.
@@
  • Phase III ($\sim$3 Myr to $\sim$50-80 Myr; the latter is the core collapse time depending on $\omega_{0}$): larger initial rotation shortens the gravothermal collapse timescale, leading to higher central density and faster mass segregation. Post-phase II angular-momentum diagnostics (e.g. $L_{\mathrm{hCCSNe}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$) rise to a peak at $\sim$10-20 Myr in the (No)K1.2 and (No)K1.8 runs then decline. The (No)K1.8 models eject many stars.
@@
  • Phase IV (core collapse to 500 Myr, the simulations’ end): post-collapse evolution is driven by binary heating in the core, producing self-similar Lagrangian-radius expansion and overall cluster expansion; models with larger $\omega_{0}$ collapse earlier and expand more. The system becomes fully mass segregated. Angular momentum is redistributed from high-mass to low-mass remnants. Low-mass BHs migrate to wider orbits. A compact central IMBH subsystem forms.
@@
We find a dependence of merger rates for stars and compact objects on the initial rotation of Pop III clusters (Fig. 9), particularly visible in the K models. The largest initial rotation ($\omega_{0}=1.8$) corresponds to the largest number of coalescences and collisions (Table 4), including the most BHBH collisions. Concerning the BHBH collisions, large initial rotation does also seem to precipitate these, with the largest number of collisions found (No)K1.8 simulations.

### 13 Apr 2026 v2

@@ 4.3.1 Structural parameter evolution
  • Phase IV: From core-collapse to 500 Myr (the simulations’ end). The Lagrangian radii evolve self-similarly, as shown in 8, driven by binary heating in the core which causes global expansion. Models with larger $\omega_{0}$ expand more by 500 Myr because they undergo earlier core collapse; their subsequent $r_{\mathrm{Lagr}}$ evolution is otherwise similar. The clusters become fully mass-segregated ($M_{\mathrm{av}}$ $\approx 200\penalty 10000\ \mathrm{M}_{\odot}{}$ centrally, $\approx 30\penalty 10000\ \mathrm{M}_{\odot}{}$ in the halo). The same long-term trends appear in Fig. 3; longer runs may amplify differences. These results are consistent with energy-flow theorems across the half-mass radius $r_{\mathrm{h}}$ (Hénon, 1975; Breen & Heggie, 2013).
@@
The gravothermal-gravogyro catastrophe is therefore present in all our Pop III cluster models, but its timescale depends strongly on the initial rotation. The GRKs have little effect on the global structural evolution.
@@ 4.3.2 Angular momentum transport
We explored whether and how angular-momentum redistribution depends on stellar evolution, GRKs, and initial bulk rotation. We tracked particles from each group over the full evolution and computed their total angular momentum squared $L_{\mathrm{group}}^{2}$ as Kamlah et al. (2022b) normalized by $L^{2}_{\omega_{0}=0.6,t=0}$ for the (No)K0.6 model. Figure 11 shows these quantities together with the group masses, $M_{\mathrm{group}}$.
@@
Phase II: $\sim$2-3 Myr. Rapid stellar evolution drives a sharp fall in $M_{\mathrm{group}}$, with higher-mass groups declining earlier owing to their shorter evolution timescales. The PISNe remove both mass and a large fraction of angular momentum: $M_{\mathrm{PISNe}}$ carried the most angular momentum before these events and is effectively depleted afterwards, causing a large loss of rotational kinetic energy.
Phase III: $\sim$3 Myr to core collapse (around $\sim$50-80 Myr depending on $\omega_{0}$. See Sect. 4.3.1 and Fig. 8). The quantity $L_{\mathrm{hCCSNe}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$ initially falls after phase II but, in the (No)K1.2 and (No)K1.8 models, it instead rises sharply and peaks at $\sim$10-20 Myr. After the peak, the quantity of (No)K1.2 shows a gradual decline, while that of (No)K1.8 drops rapidly. Other mass groups follow the same pattern, except for $M_{\mathrm{PISNe}}$. In (No)K1.8 the quantity attains a local minimum at core collapse, when the cluster is essentially fully mass-segregated (Fig. 8).
@@
Compared to rotating Pop II clusters (Kamlah et al., 2022b), the detailed transport history differs because of the IMF and stellar evolution, but the gravothermal-gravogyro catastrophe remains visible in both cases. The GRKs again produce no strong effect.
@@ 4.4 Collision and coalescence events
We defined a ‘coalescence’ as the merger of a nearly circular binary, while ‘collisions’ are a direct physical impact either on a hyperbolic or highly eccentric orbit; this is also how the Nbody6++GPU code separates them. Under our initial conditions and Pop III stellar evolution prescriptions, we find collisions of BHBHs, MSBHs, and MSMSs, while coalescences may involve MS, HG, CHeB, ShHeB, HeMS, NS, and BH progenitors. Figure 12 shows the product masses, $m_{\mathrm{coal}}$ and $m_{\mathrm{coll}}$, versus $r_{\mathrm{dens}}$. For coalescence, only MS, CHeB, ShHeB, and BH products appear, consistent with the merger-product treatment of Hurley et al. (2002) and Kamlah et al. (2022a). Event totals are given in Table 4.
@@
Our suite contains one realization per set-up ($N=1$ per model ID). For integer event counts, $N$ (specific coalescence and collision types, compact-object merger channels), the counting uncertainty is well approximated by Poisson statistics, with a reference $1\sigma$ uncertainty of $\sigma_{N}\simeq\sqrt{N}$ (relative uncertainty $\sigma_{N}/N\simeq 1/\sqrt{N}$). Consequently, differences of a few events are not, by themselves, strong evidence of a systematic dependence on $\omega_{0}$ when NN is in the single-digit to few-tens regime [^8] By contrast, trends inferred from continuous structural quantities (e.g. Figs. 3 and 8, and the angular-momentum transport in Fig. 11) and from large-number counts (e.g. total escaper numbers; Table 5) are dynamically more robust in our dataset..
@@
The coalescence events show no marked spatial differences: most occur within the central 2.5 pc, i.e. inside $r_{h}$ during the first $\sim$40 Myr (Fig. 3). They are mostly completed by 4-5 Myr (pre-core-collapse; Fig. 9). Many of these early mergers populate the PISN mass gap, indicating that stellar mergers in early star cluster evolution can seed BHs in and around the gap (see e.g. Costa et al., 2022; Ballone et al., 2022, for recent hydrodynamical simulations of stellar collisions on this topic). With our flat IMF, Pop III stellar evolution, and absent wind mass loss, producing BHs above the PISN gap is easier here than in typical young massive clusters (e.g. Rizzuto et al., 2022).
@@
Most coalescences produce CHeB stars and the most massive products, typically occurring within a few million years in all runs. Faster initial rotation is consistent with more CHeB coalescences (Fig. 9), but Poisson uncertainties and covariances between channels in a single realization make this indicative. The numbers of BH, ShHeB, and MS coalescences are similar across simulations. These mergers produce stars (mainly CHeB) with masses inside and above the pair-instability gap; some IMBHs end up in the pair-instability gap after accreting stellar material and grow further.
@@
Collisions of MSMS, BHMS and BHBH offer useful diagnostics. Collision-produced IMBH masses are broadly similar across models (Fig. 12), except for NoK1.2 (IMBH masses $\approx$ 1000 $\mathrm{M}_{\odot}$), which is not seen in K1.2. Figure 9 suggests more MSMS and BHBH collisions at a higher initial rotation, but given the small absolute counts the implied trend should be read as suggestive and may be affected by stochastic fluctuations. Notably, K models show multi-generation IMBH growth.
@@
Figure 13 shows the primary versus secondary masses for MSMS, BHMS, and BHBH collisions across all runs; the distributions are similar regardless of rotation or GRKs. Contrary to theoretical expectations (e.g. Morawski et al., 2018, 2019) [^9] From these theories, we would expect that mergers have a very low GRK velocity, which would make them promising candidates for growing seed BHs for galactic nuclei., Fig. 14 demonstrates that most GRK velocities, $v_{\mathrm{GRk}}$, exceed cluster escape speeds in the core, so merger remnants are often ejected even for $q\approx 1$. Hence, IMBHs of $1000\mathrm{M}_{\odot}{}$ are unlikely to form via hierarchical merging in Pop III clusters unless additional processes retain the remnants [^10] A surrounding dark matter halo is unlikely to help to retain merger remnants. Even assuming a NFW big-halo of 100 times of the cluster total mass with 10 times of the cluster tidal radius will only contribute to a few $\mathrm{km\penalty 10000\ s}^{-1}$ of the escape speed in the core., such as strong gas inflow after cluster formation causing extreme contraction of the cluster (Kroupa et al., 2020).
@@
To summarize, with increasing initial rotation, $\omega_{0}$, coalescences occur earlier and more frequently, and there is a tentative rise in MSMS and BHBH collision counts limited by low-number statistics. Enabling GRKs generates kick velocities that frequently exceed the cluster escape speed, favouring remnant ejection and thus suppressing collision-built IMBHs, even though multi-generation IMBH growth can still occur.
@@ 4.5 Escaper stars
Figure 4: Escaping single stars in eight simulations: cumulative counts and masses. Top panel: Cumulative number of escaping single stars, $n_{\mathrm{esc,singles}}$. Second panel: Total mass in escaping single stars, $M_{\mathrm{esc,singles}}$. Remaining panels: Cumulative counts of escaping single MS, CHeB, ShHeB, NS, and BH stars, $n_{\mathrm{escMS,s}}$, $n_{\mathrm{escCHeB,s}}$, $n_{\mathrm{escShHeB,s}}$, $n_{\mathrm{escNS,s}}$, and $n_{\mathrm{escBH,s}}$. Solid curves show K models and dash-dotted curves show NoK models.
Figure 4 (single escapers), Fig. 10 (binary escapers), and Table 5 show that a higher initial bulk rotation, $\omega_{0}$, is associated with earlier and stronger tidal mass loss in both NoK and K models, most clearly for (No)K1.8 and (No)K1.2. The (No)K0.6 models show a non-monotonic behaviour, having $n_{\mathrm{esc,singles}}$, $n_{\mathrm{esc,binarymembers}}$ and $M_{\mathrm{esc,singles}}$ comparable to or below the non-rotating case. A plausible physical interpretation is that two effects compete: (i) rotating King initial conditions are more compact than for their non-rotating counterparts (Sect. 1), which can reduce early tidal overflow, while (ii) stronger rotation accelerates early angular-momentum transport and structural readjustment (Sect. 4.3.1), which enhances mass loss once the system expands and interacts with the tidal boundary. In this framework, $\omega_{0}=0.6$ can fall in an intermediate regime where the increased initial compactness dominates the net mass loss, whereas for larger $\omega_{0}$ the rotation-driven early evolution dominates. This tidal field mass loss modifies cluster timescales and links directly to the structural and angular-momentum evolution discussed in Sects. 4.3.1 and 4.3.2.
@@
We next considered evolved phases, starting with escaping single CHeB stars ($n_{\mathrm{escCHeB,s}}$; Fig. 4 panel four). Their behaviour resembles that of escaping MSs, though differences between (No)K1.2 and (No)K1.8 are reduced. Many escaping MSs rapidly become CHeBs, and likewise many CHeBs shortly become ShHeBs (see Fig. 5). After formation, most will remain in the ShHeB phase for a comparatively long time.
@@
For escaping binary components excluding compact objects, a few MS escapers (counted by $n_{\mathrm{escMS,b}}$) occur only in the (No)K1.2 and (No)K1.8 runs; they are mostly from primordial binaries, low-mass, and unlikely to form IMBHs. Binary CHeB escapers ($n_{\mathrm{escBH,b}}$) are rarer still. Binary ShHeB losses ($n_{\mathrm{escShHeB,b}}$) are roughly twice as large and occur earlier in the $\omega_{0}=1.2,1.8$ models than in the $\omega_{0}=0.0,0.6$ models. Many late-time escaping BHBH binaries contain two IMBHs, and thus cannot fuel hierarchical in-cluster growth. The stronger-rotation models are therefore consistent with a reduced retained BH progenitor and seed population at late times.
Overall, enabling GRKs produces the charted differences: with GRKs enabled the (No)K1.8 models show more CHeB single-star losses and fewer MS binary losses, while the (No)K1.2 models show the opposite trend. We again note that this is limited by small-number statistics.
@@ 5 Summary and conclusions
We present a set of a few pilot direct N-body simulations that probe how bulk rotation and ultra-low metallicity with a top-heavy IMF shape Pop III cluster dynamics. Rotation follows generalized rotating King models (Einsel & Spurzem, 1999) with rigid core rotation and differentially decreasing near the half-mass radius and into the halo. Kamlah et al. (2022b) analysed the general evolution and angular momentum transport in such models of Pop II star clusters; we find generally analogous dynamics, such as the gravogyro and gravothermal collapse, enhanced by the top-heavy IMF (8.08.0300.0300.0 $\mathrm{M}_{\odot}$, Lazar & Bromm, 2022). We study the impact of switching on and off GRKs (Arca Sedda et al. 2024), and how the newly implemented stellar evolution for extremely metal-poor (EMP) stars (here used $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, Tanikawa et al., 2020) affects the formation and evolution of compact objects. While the mass function of Pop III star clusters is highly uncertain (see Sect. 1 and Stacy et al., 2016; Liu et al., 2021a), our pilot simulations model very massive Pop III star clusters (Table 1) to cover an unexplored parameter range and also because of stronger and faster dynamical effects, such as massive object formation, at high density (Vergara et al., 2025). Direct N-body simulations using the PeTar code (Wang et al., 2020) for Pop III star clusters (Liu et al., 2024a, b, 2025) have used smaller cluster masses and particle numbers, requiring many independent realizations to average $N$ fluctuations. On the contrary, we used very large particle numbers ($1.01\times 10^{5}$ initially), which suffice to probe global dynamical effects of rotation and gravogyro-gravothermal collapse, but still provide a relatively small database for rare binaries and escapers (e.g. those involving BHs). We ran the models for 500 Myr, using one non-rotating King model and three rotating ones (Einsel & Spurzem 1999, $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$), each with and without GR merger recoil kicks.
@@
Primordial binary statistics are another major uncertainty for Pop III clusters. In this work we fixed the initial binary fraction to $f_{\mathrm{b}}=1\%$ (Table 1), which should be regarded as a low-$f_{\mathrm{b}}$ limiting case rather than a best-estimate value. This choice suppresses IMBH formation pathways tied to primordial binary evolution (in particular Type-I and Type-III; Sect. 4.1) and, by construction, increases the relative importance of dynamical formation channels in our event budget. If $f_{\mathrm{b}}$ were higher, we would expect (i) a larger contribution of mergers from primordial binary channel, and (ii) stronger binary heating, which can delay gravothermal collapse and modify the balance between in-cluster hierarchical growth and dynamical ejections. Predicting the net change in the total IMBH production efficiency is therefore not straightforward without additional simulations, but the dominance of dynamical channels found here should be interpreted as an upper bound. A dedicated scan in $f_{\mathrm{b}}$ is left for future work.
@@
The global dynamical evolution can be divided into four phases (Figs. 3, 8, and 11):
@@
  • Phase I (0-2 Myr): Global mass ($M_{\mathrm{cl}}$) and tidal radius ($r_{\mathrm{tid}}$) remain essentially constant because mass loss is negligible; however, initial bulk rotation ($\omega_{0}$) controls the structural response. Rapid angular-momentum transport occurs earlier and more strongly for larger $\omega_{0}$, producing a brief, synchronous bump in the Lagrangian radii near $\sim$0.1 Myr, inner-radius contraction and outer-radius expansion, and temporary rises then falls in $r_{\mathrm{h}}$, $r_{\mathrm{c}}$, and $m_{\mathrm{c}}$ in the faster-rotating models. Low and zero rotation models show little change in $r_{\mathrm{h}}$ over this phase. Also, $M_{\mathrm{group}}$ evolves similarly across mass groups, while $L_{\mathrm{group}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$ is larger for higher $\omega_{0}$, reflecting early outward angular-momentum migration.
@@
  • Phase II ($\sim$2-3 Myr): Rapid stellar-evolution mass loss (notably PISNe and compact-object formation) causes a sharp drop in total mass and $r_{\mathrm{tid}}$, and produces a rapid increase in $r_{\mathrm{h}}$ and $r_{\mathrm{c}}$. This is due to PISNe both remove mass and carry away a large fraction of the total angular momentum, while escapers (natal or GR kicks) and dynamical encounters add to mass loss and alter kinematics through fallback-dependent kicks. Measures of central segregation ($M_{\mathrm{av}}$ within small $r_{\mathrm{Lagr}}$) show a similar qualitative behaviour across $\omega_{0}$ at this stage.
@@
  • Phase III ($\sim$3 Myr to $\sim$50-80 Myr; the latter is the core collapse time depending on $\omega_{0}$): Larger initial rotation shortens the gravothermal collapse timescale, leading to a higher central density and faster mass segregation. Post-phase II angular-momentum diagnostics (e.g. $L_{\mathrm{hCCSNe}}^{2}/L^{2}_{\omega_{0}=0.6,t=0}$) rise to a peak at $\sim$10-20 Myr in the (No)K1.2 and (No)K1.8 runs then decline. The (No)K1.8 models eject many stars.
@@
  • Phase IV (core collapse to 500 Myr, the simulations’ end): Post-collapse evolution is driven by binary heating in the core, producing self-similar Lagrangian-radius expansion and overall cluster expansion; models with larger $\omega_{0}$ collapse earlier and expand more. The system becomes fully mass-segregated. Angular momentum is redistributed from high-mass to low-mass remnants. Low-mass BHs migrate to wider orbits. A compact central IMBH subsystem forms.
@@
We find a dependence of merger rates for stars and compact objects on the initial rotation of Pop III clusters (Fig. 9), particularly visible in the K models. The largest initial rotation ($\omega_{0}=1.8$) corresponds to the largest number of coalescences and collisions (Table 4), including the most BHBH collisions. Concerning the BHBH collisions, large initial rotation does also seem to precipitate these, with the largest number of collisions found in (No)K1.8 simulations.